# Are Fact-Checking Tools Helpful? An Evaluation of Google Fact Check

### v3

@@ II-E Input Claim Variation
To collect the varied claims addressing the same issue, we involved the coders who contributed in Sections II-B and II-C. Each coder was assigned a datasheet copy listing the fact-checking results rated “relevant” in Section II-B1. The coders first individually reviewed and tagged each input claim with one or more keywords that can summarize the key topic of the claim. Then, the claims with identical or similar tags were grouped by manually dragging the corresponding rows in the datasheet. We considered “similar” tags as well because coders may forget what tags they previously used when coding for more than 200 claims. For instance, if there is a claim “There is a relationship between COVID-19 and the 5G network” at the beginning of the datasheet, it could be tagged “COVID-19 and 5G”; after a long coding process, the coder may see another similar claim “Coronavirus cases linked to 5G rollout” at the end of the datasheet, but they may forget the tag attached to the previous claim and thus use a new tag “Coronavirus and 5G”, although these two claims should share the same tag and grouped because they essentially described the same story. Therefore, after the initial tagging, each coder reviewed, grouped, and merged any tags that essentially meant the same. Then, we compared the coding results from the three coders and had an open discussion to resolve any disagreements. Note that the goal of tagging is to find out the claims addressing the same issue yet described differently, so it is acceptable if the coders tagged claims with different keywords as long as they grouped the same claims. For instance, the two claims mentioned above could be tagged as “COVID-19 and 5G” by one coder and “Coronavirus and 5G” by another, but it is acceptable if they both found and grouped these two claims addressing the same issue.
@@
$J(S_{1},S_{2})=\frac{|S_{1}\cap S_{2}|}{|S_{1}\cup S_{2}|}=\frac{\text{Number of Overlapped Results}}{\text{Number of Unique Results in Total}}$ (1)
Since the Jaccard index is normally used to calculate the congruence between two groups, we calculated the average Jaccard index of all the permutation pairs for the tags involving more than two input claims. For instance, if there are three claims, A, B, and C, with the same tag, we first calculated the Jaccard indices of A and B, A and C, and B and C, respectively, then calculated the average. The results are shown in Section III-E.
III Results and Discussion
III-A Statistics of Fact-Checking Results
Among the 1,000 false claims we leveraged, 842 (84.2%) failed to get any results, 101 (10.1%) received one result each, and 57 (5.7%) received two or more results. The total number of the results is 290. Table I shows more detail. We also found that all the reviewed claims are unique, i.e., no sources reviewed the identical claims. This corresponds to the findings in the previous studies [11, 12].
TABLE I: The number of fact-checking results received by a claim and the corresponding number of input claims involved. 158 (15.8%) input claims obtained 290 fact-checking results in total from Google Fact Check.
@@ III-A Statistics of Fact-Checking Results
As for the relevance of results, one result was not in English and thus was ignored; 273 (94.46%) of the remaining 289 results were rated relevant to the input claims by three coders because their claims assessed by fact-checkers were in English and focused on the same topic as the corresponding input claims; the remaining 16 (5.54%) were rated irrelevant because they deviated from the main topic of the input claims. Table II presents the results.
TABLE II: The distribution of the relevant and irrelevant results rated by three coders. Note that we excluded one non-English result from the 290 results returned by Google Fact Check, so the total number of valid results was 289.
@@
Even though most (94.46%) of the results were rated relevant to the input claims, the large portion of claims without results limits the usefulness of collecting fact-checking information. According to the About page of Google Fact Check, the results were provided spontaneously by fact-checkers, i.e., when publishing a fact-checking report, publishers can opt to attach a ClaimReview markup to make it detectable by a search engine [8]. Therefore, a possible explanation is that more fact-checking results did exist somewhere online but were not detectable by Google Fact Check, of which the usability was thus limited.
III-B Statistics of Fact-Checking Verdicts
To analyze the fact-checking verdicts, we only considered the 273 fact-checking results rated relevant to the original input claims in the section above. We further excluded a result in which the fact-checking verdict was not in English, so the total number of results to analyze was reduced to 272. Table III shows the distribution of fact-checking verdicts. 217 (79.78%) verdicts were “False” and 32 (11.76%) were “Partly False”. These two categories debunking inaccuracies with clear evidence accounted for 91.54% in total. As for the remaining, 22 (8.09%) were “Unratable” without conclusive evidence to prove true or false, and only one (0.37%) was “True”. Since the claims leveraged were from the FakeCovid dataset and were manually scrutinized by editors, they are more likely to be problematic, influential, and thus worth being fact-checked. This may explain why most of the verdicts we obtained were negative. Even so, it is reasonable to assume that professionals have a higher sensitivity to hot topics and misinformation so that these fact-checked claims are likely to reflect public focus in general.
TABLE III: The distribution of fact-checking verdicts. We excluded 16 irrelevant results and one result whose fact-checking verdict was not in English, so the total number of results involved was reduced to 272.
@@ III-B Statistics of Fact-Checking Verdicts
III-C Statistics of Source Reliability
Table IV lists the distribution of the fact-checking sources referenced by the 272 valid fact-checking results in the section above and their reliabilities and political leanings. The top five sources referenced the most are PolitiFact (n = 53, 19.49%), AFP Fact Check (n = 47, 17.28%), Full Fact (n = 26, 9.56%), Health Feedback (n = 20, 7.35%), and FactCheck (n = 16, 5.88%). In total, they were referenced by 162 (59.56%) results. 11 (4.04%) results were returned without source information.
We referred to two websites, Interactive Media Bias Chart and Media Bias/Fact Check, to evaluate each source’s reliability and political leaning. Eight (32%) of the 25 sources involved were rated by both websites, six (24%) by one only, and 11 (44%) by neither one. Among the sources rated by at least one website (n = 14, 56%), all were rated “Trustworthy” or “Relatively Trustworthy”, indicating high reliability of their fact-checking verdicts; as for political leaning, 14 (63.64%) of the 22 ratings in the table are “Center”, indicating a relatively unbiased stance in general.
TABLE IV: The distribution of the fact-checking sources referenced by 272 valid fact-checking results.
@@ III-C Statistics of Source Reliability
Alt News 7 2.57% n/a Trustworthy n/a Left
@@
Ghana Fact 4 1.47% n/a n/a n/a n/a
Australian Associated Press 4 1.47% n/a Trustworthy n/a Center
@@
Check4Spam 3 1.10% n/a n/a n/a n/a
Newsmeter 3 1.10% n/a n/a n/a n/a
Newsmobile 2 0.74% n/a n/a n/a n/a
Newschecker 2 0.74% n/a n/a n/a n/a
The Quint 2 0.74% n/a Trustworthy n/a Left
THIP Media 2 0.74% n/a n/a n/a n/a
The Journal 2 0.74% n/a n/a n/a n/a
FactCheckHub 1 0.37% n/a n/a n/a n/a
Namibia Fact Check 1 0.37% n/a n/a n/a n/a
Africa Check 1 0.37% n/a Trustworthy n/a Center
FactRakers 1 0.37% n/a n/a n/a n/a
(null) 11 4.04% n/a n/a n/a n/a
III-D Relationships Between Claims and Results
III-D1 Number of Results vs. Characteristics of Input Claims
Table V lists Spearman’s rank correlation coefficients between the characteristics of input claims and their corresponding numbers of fact-checking results. No significant correlation was detected as no coefficient was greater than 0.4 or smaller than -0.4, indicating a weak correlation [22].
TABLE V: The Spearman’s rank correlation coefficients between the characteristics of the input claims and the number of fact-checking results.
@@ III-D Relationships Between Claims and Results
III-D2 Relevance of Results vs. Characteristics of Input Claims
We performed the Kruskal-Wallis H test to detect the relationship between the relevance of results and the characteristics of input claims. Table VI shows the results. There was no p-value less than 0.05, although the p-value for emotional tone (p = 0.051) was close to the critical value when we accepted tie raking. Therefore, no significant difference in the characteristics of input claims was observed.
TABLE VI: Kruskal-Wallis H test results for the characteristics of input claims and result relevance.
@@
III-D3 Fact-Checking Verdicts vs. Characteristics of Input Claims
We performed the Kruskal-Wallis H test between the characteristics of input claims and the fact-checking verdicts. As Table VII shows, most of the p-values were greater than 0.05, indicating insignificant difference in the characteristics of input claims. The only significant result was observed for emotional tone (p = 0.014) when tie ranking was adopted.
TABLE VII: Kruskal-Wallis H test results for the characteristics of input claims and fact-checking verdicts.
@@
III-D4 Fact-Checking Sources vs. Characteristics of Input Claims
We performed the Kruskal-Wallis H test for the characteristics of input claims based on fact-checking sources. The results are shown in Table VIII. We observed significant results for word count (p = 0.035 without ties and p = 0.034 with ties) and dictionary words (p = 0.001 without ties and p = 0 with ties), indicating that different sources tend to fact-check claims with varied lengths and using dictionary words at different extents. All the other p-values were greater than 0.05 and thus insignificant.
TABLE VIII: Kruskal-Wallis H test results for the characteristics of input claims and fact-checking sources.
@@
III-E Input Claim Variation
We calculated the Jaccard index to quantify the congruence of the fact-checking results regarding claim variations, i.e., claims addressing the same issue but in different descriptive ways. Table IX shows the numbers of fact-checking results in union and intersection sets and the corresponding Jaccard indices regarding the 21 tags involving at least two claim variations. We calculated the regular Jaccard index for the 18 (85.71%) tags involving two input claim variations and the averaged Jaccard index for the remaining three (14.29%) tags involving three or more claim variations. 17 (80.95%) tags achieved low similarities (not greater than 0.5). Even though three (14.29%) tags achieved perfect similarity (100%), they each only involved two claim variations that received one identical fact-checking result. Therefore, input claims in different descriptions are not likely to obtain identical fact-checking results. Even though descriptive variations tend to cover nuanced details that may result in different results, such a situation is relatable to real life: when people are concerned about the same issue, they may individually search for related information in their own words for verification. Therefore, our investigation regarding such an issue is insightful.
TABLE IX: The Jaccard Index of each topic.
@@ IV Conclusion and Recommendations
In this study, we evaluated the performance of Google Fact Check, a Google-based search engine for fact-checking results. Our study was motivated by two problems. First, even though Google Fact Check brings great convenience to fact-checking tasks and has been leveraged as a core component of some automated fact-checking frameworks, its performance should be fully investigated before rolling out, such as whether the fact-checking results it returned are helpful and reliable. Second, it was unclear whether different fact-checking sources tend to fact-check the same claims and whether their verdicts tend to be congruent, and a search engine collecting fact-checking results from various sources could pave the way for an investigation. We obtained the fact-checking results for 1,000 claims about COVID-19 via Google Fact Check API and analyzed them from the perspectives of result quality and the relationships between input claims and fact-checking results. We found that most claims did not obtain fact-checking results, even though the returned results were relatively relevant to the input claims and reliable, and multiple sources are not likely to check exactly identical claims. Besides, we did not detect significant correlations between the linguistic characteristics of input claims and the results they obtained, except that fact-checking sources tend to check the claims with different lengths and the usage of dictionary words. We also found that the variations of input claim wording are likely to obtain different fact-checking results. Based on these findings, we suggested that the usability of Google Fact Check can be improved by optimizing the algorithms and enhancing collaboration with fact-checking sources to compile more fact-checking results available online. Users may slightly adjust their wording to collect more potential fact-checking results, and the linguistic characteristics of input claims are not likely to significantly influence the quality of results.
@@
There are some limitations that we could not bypass in this study. First, due to the large amount of input claims we leveraged (n = 1,000), we obtained the fact-checking results via API programming. However, we noticed in our pre-study experiment that the results from the API could be slightly different from the results page. We did not delve into it because API programming is common in software development, such as automated fact-checking frameworks. Considering that our results were compiled in December 2022, future work may continue to investigate whether such a difference should be a big concern and whether the performance of Google Fact Check has been significantly optimized. Second, we noticed that a small number of claims seemed irrelevant to COVID-19 (e.g., “Trump and McConnell are blocking stimulus checks for Americans married to immigrants”). However, we then found them still in the context of COVID-19. With the large amount of claims we leveraged (n = 1,000), they are less likely to compromise the validity of the study. That said, future work can leverage the claims more directly related to a topic. Finally, the subjectivity in this study was unavoidable. Even though we recruited three coders, we could not guarantee the accuracy of the coding results. For instance, in Section III-B, there was a “True” verdict, which should have been rated “irrelevant” in Section II-B1 and excluded from the analyses because its claim assessed by fact-checkers (i.e., “Disposable masks should always be worn colored-side-out”) was not addressing the same issue as the original input claim (i.e., “Two ways of wearing a surgical mask: Colored side out if you are sick and white side out if you do not want to become sick”), that is, the verdict “True” did not indicate that the input claim was true but rather another irrelevant claim, but it was rated “relevant” by two of the three coders and still analyzed thereafter. As for the evaluation of source reliability, even though we referred to two professional websites for media evaluation, as emphasized on their About pages, their ratings could also be more or less biased. Even though subjectivity is unavoidable in any qualitative research, future work may recruit more coders and leverage more reputable rating sources and criteria to further reduce the potential influence of subjectivity.

### v5

@@ 2.5 Input Claim Variation
To collect the varied claims addressing the same issue, we involved the coders who contributed in Sections 2.2 and 2.3. Each coder was assigned a datasheet copy listing the fact-checking results rated “relevant” in Section 2.2.1. The coders first individually reviewed and tagged each input claim with one or more keywords that can summarize the key topic of the claim. Then, the claims with identical or similar tags were grouped by manually dragging the corresponding rows in the datasheet. We considered “similar” tags as well because coders may forget what tags they previously used when coding for more than 200 claims. For instance, if there is a claim “There is a relationship between COVID-19 and the 5G network” at the beginning of the datasheet, it could be tagged “COVID-19 and 5G”; after a long coding process, the coder may see another similar claim “Coronavirus cases linked to 5G rollout” at the end of the datasheet, but they may forget the tag attached to the previous claim and thus use a new tag “Coronavirus and 5G,” although these two claims should share the same tag and grouped because they essentially described the same story. Therefore, after the initial tagging, each coder reviewed, grouped, and merged any tags that essentially meant the same. Then, we compared the coding results from the three coders and had an open discussion to resolve any disagreements. Note that the goal of tagging is to find out the claims addressing the same issue yet described differently, so it is acceptable if the coders tagged claims with different keywords as long as they grouped the same claims. For instance, the two claims mentioned above could be tagged as “COVID-19 and 5G” by one coder and “Coronavirus and 5G” by another, but it is acceptable if they both found and grouped these two claims addressing the same issue.
@@
$J(S_{1},S_{2})=\frac{|S_{1}\cap S_{2}|}{|S_{1}\cup S_{2}|}=\frac{\text{Number of Results in the Intersection}}{\text{Number of Results in the Union}}$ (1)
Since the Jaccard index is normally used to calculate the congruence between two groups, we calculated the average Jaccard index of all the permutation pairs for the tags involving more than two input claims. For instance, if there are three claims, A, B, and C, with the same tag, we first calculated the Jaccard indices of A and B, A and C, and B and C, respectively, then calculated the average. The results are shown in Section 3.5.
3 Results
3.1 Statistics of Fact-Checking Results
Among the 1,000 false claims we leveraged, 842 (84.2%) failed to get any results, 101 (10.1%) received one result each, and 57 (5.7%) received two or more results. The total number of the results is 290. Table 1 shows more detail. We also found that all the reviewed claims are unique, i.e., no sources reviewed the identical claims. This corresponds to the findings in the previous studies [17, 19].
Table 1: The number of fact-checking results received by a claim and the corresponding number of input claims involved. 158 (15.8%) input claims obtained 290 fact-checking results in total from Google Fact Check.
@@ 3.1 Statistics of Fact-Checking Results
As for the relevance of results, one result was not in English and thus was ignored; 273 (94.46%) of the remaining 289 results were rated relevant to the input claims by three coders because their claims assessed by fact-checkers were in English and focused on the same topic as the corresponding input claims; the remaining 16 (5.54%) were rated irrelevant because they deviated from the main topic of the input claims. Table 2 presents the results.
Table 2: The distribution of the relevant and irrelevant results rated by three coders. Note that we excluded one non-English result from the 290 results returned by Google Fact Check, so the total number of valid results was 289.
@@
Even though most (94.46%) of the results were rated relevant to the input claims, the lack of results regarding 84.2% of the input claims may not well help users fact-check the issues they have concerns about. According to the About page of Google Fact Check, the results were provided spontaneously by fact-checkers, i.e., when publishing a fact-checking report, publishers can opt to attach a ClaimReview markup to make it detectable by a search engine [3]. Therefore, a possible explanation is that more fact-checking results did exist somewhere online but were not detectable by Google Fact Check, of which the usability was thus limited.
3.2 Statistics of Fact-Checking Verdicts
To analyze the fact-checking verdicts, we only considered the 273 fact-checking results rated relevant to the original input claims in the section above. We further excluded a result in which the fact-checking verdict was not in English, so the total number of results to analyze was reduced to 272. Table 3 shows the distribution of fact-checking verdicts. 217 (79.78%) verdicts were “False” and 32 (11.76%) were “Partly False.” These two categories debunking inaccuracies with clear evidence accounted for 91.54% in total. As for the remaining, 22 (8.09%) were “Unratable” without conclusive evidence to prove true or false, and only one (0.37%) was “True.” Since the claims leveraged were from the FakeCovid dataset and were manually scrutinized by editors, they are more likely to be problematic, influential, and thus worth being fact-checked. This may explain why most of the verdicts we obtained were negative. Even so, it is reasonable to assume that professionals have a higher sensitivity to hot topics and misinformation so that these fact-checked claims are likely to reflect public focus in general.
Table 3: The distribution of fact-checking verdicts. We excluded 16 irrelevant results and one result whose fact-checking verdict was not in English, so the total number of results involved was reduced to 272.
@@ 3.2 Statistics of Fact-Checking Verdicts
3.3 Statistics of Source Reliability
Table 4 lists the distribution of the fact-checking sources referenced by the 272 valid fact-checking results in the section above and their reliabilities and political leanings. The top five sources referenced the most are PolitiFact (n = 53, 19.49%), AFP Fact Check (n = 47, 17.28%), Full Fact (n = 26, 9.56%), Health Feedback (n = 20, 7.35%), and FactCheck (n = 16, 5.88%). In total, they were referenced by 162 (59.56%) results. 11 (4.04%) results were returned without source information.
We referred to two websites, Interactive Media Bias Chart and Media Bias/Fact Check, to evaluate each source’s reliability and political leaning. Eight (32%) of the 25 sources involved were rated by both websites, six (24%) by one only, and 11 (44%) by neither one. Among the sources rated by at least one website (n = 14, 56%), all were rated “Trustworthy” or “Relatively Trustworthy,” indicating high reliability of their fact-checking verdicts; as for political leaning, 14 (63.64%) of the 22 ratings in the table are “Center,” indicating a relatively unbiased stance in general.
Table 4: The distribution of the fact-checking sources referenced by 272 valid fact-checking results.
@@ 3.3 Statistics of Source Reliability
Alt News 7 2.57% Trustworthy Left
@@
Ghana Fact 4 1.47%
Australian Associated Press 4 1.47% Trustworthy Center
@@
Check4Spam 3 1.10%
Newsmeter 3 1.10%
Newsmobile 2 0.74%
Newschecker 2 0.74%
The Quint 2 0.74% Trustworthy Left
THIP Media 2 0.74%
The Journal 2 0.74%
FactCheckHub 1 0.37%
Namibia Fact Check 1 0.37%
Africa Check 1 0.37% Trustworthy Center
FactRakers 1 0.37%
(null) 11 4.04%
3.4 Correlation Between Claims and Results
3.4.1 Number of Results vs. Characteristics of Input Claims
Table 5 lists Spearman’s rank correlation coefficients between the characteristics of input claims and their corresponding numbers of fact-checking results. No significant correlation was detected as no coefficient was greater than 0.4 or smaller than -0.4, indicating a weak correlation [8].
Table 5: The Spearman’s rank correlation coefficients between the characteristics of the input claims and the number of fact-checking results.
@@ 3.4 Correlation Between Claims and Results
3.4.2 Relevance of Results vs. Characteristics of Input Claims
We performed the Kruskal-Wallis H test to detect the relationship between the relevance of results and the characteristics of input claims. Table 6 shows the results. There was no p-value less than 0.05, although the p-value for emotional tone (p = 0.051) was close to the critical value when we accepted tie raking. Therefore, no significant difference in the characteristics of input claims was observed.
Table 6: Kruskal-Wallis H test results for the characteristics of input claims and result relevance.
@@
3.4.3 Fact-Checking Verdicts vs. Characteristics of Input Claims
We performed the Kruskal-Wallis H test between the characteristics of input claims and the fact-checking verdicts. As Table 7 shows, most of the p-values were greater than 0.05, indicating insignificant difference in the characteristics of input claims. The only significant result was observed for emotional tone (p = 0.014) when tie ranking was adopted.
Table 7: Kruskal-Wallis H test results for the characteristics of input claims and fact-checking verdicts.
@@
3.4.4 Fact-Checking Sources vs. Characteristics of Input Claims
We performed the Kruskal-Wallis H test for the characteristics of input claims based on fact-checking sources. The results are shown in Table 8. We observed significant results for word count (p = 0.035 without ties and p = 0.034 with ties) and dictionary words (p = 0.001 without ties and p = 0 with ties), indicating that different sources tend to fact-check claims with varied lengths and using dictionary words at different extents. All the other p-values were greater than 0.05 and thus insignificant.
Table 8: Kruskal-Wallis H test results for the characteristics of input claims and fact-checking sources.
@@
3.5 Input Claim Variation
We calculated the Jaccard index to quantify the congruence of the fact-checking results regarding claim variations, i.e., claims addressing the same issue but in different descriptive ways. Table 9 shows the numbers of fact-checking results in union and intersection sets and the corresponding Jaccard indices regarding the 21 tags involving at least two claim variations. We calculated the regular Jaccard index for the 18 (85.71%) tags involving two input claim variations and the averaged Jaccard index for the remaining three (14.29%) tags involving three or more claim variations. 17 (80.95%) tags achieved low similarities (not greater than 0.5). Even though three (14.29%) tags achieved perfect similarity (100%), they each only involved two claim variations that received one identical fact-checking result. Therefore, input claims in different descriptions are not likely to obtain identical fact-checking results. Even though descriptive variations tend to cover nuanced details that may produce different results, such a situation is relatable to real life: when people are concerned about the same issue, they may individually search for related information in their own words for verification. Therefore, our investigation regarding such an issue is insightful.
Table 9: The Jaccard Index of each topic.
@@ 4 Discussion and Conclusion
In this study, we explored the promise of fact-checking-specific search engines by evaluating the performance of Google Fact Check, a Google-based search engine for fact-checking results and the only fact-checking-specific search engine as of this study. Our study was motivated by two practical issues. First, even though Google Fact Check and other similar tools bring great convenience to fact-checking tasks and have been leveraged as a core component of some automated fact-checking frameworks, it is necessary to validate its usability, such as whether the fact-checking results it renders are helpful and reliable. Second, it was unclear whether different fact-checking sources tend to fact-check the same claims and whether their verdicts tend to be congruent, and a search engine collecting fact-checking results from various sources could pave the way for an investigation. We retrieved the fact-checking results for 1,000 COVID-19-related false claims via Google Fact Check API and analyzed them from the perspectives of result quality and the correlations between input claims and fact-checking results. We found that most claims did not retrieve fact-checking results, even though the returned results were relatively relevant to the corresponding input claims and tended to be reliable. Furthermore, we did not detect significant correlations between the linguistic characteristics of input claims and the corresponding fact-checking results, except that different fact-checking sources are less likely to repeatedly fact-check identical content and tend to check the claims with various lengths and the usage of dictionary words. We also found that the variations of input claim wording are likely to result in different fact-checking information. Based on these findings, we suggest that the quantities of the fact-checking results rendered by Google Fact Check and similar tools can be optimized by, for instance, enhancing collaborations with fact-checking sources to broaden the result scope. Users may not necessarily worry about the discrepancies among sources regarding the same content, which is not likely to be repeatedly checked across sources; however, they could slightly adjust the input wording for more potential fact-checking results, although the linguistic characteristics of input claims are not likely to significantly influence the quality of results.
@@
There are some limitations that we could not bypass in this study. First, due to the large amount of input claims we leveraged (n = 1,000), we retrieved the fact-checking results via API programming. However, we noticed in our preliminary experiment that the results from the API may not be exactly the same as those on the results page. We did not delve into this observation because API programming is prevalent in software development, such as automated fact-checking frameworks. Considering that our results were compiled in December 2022, future work may continue to investigate whether such a difference should be a big concern and whether the performance of Google Fact Check has been significantly optimized since then. Second, we noticed that a small number of claims seem irrelevant to COVID-19 (e.g., “Trump and McConnell are blocking stimulus checks for Americans married to immigrants”). However, we then found them still in the context of COVID-19. With the large amount of claims we leveraged (n = 1,000), they are less likely to compromise the validity of the study. That said, future work may leverage the claims more directly related to a topic. Third, the subjectivity in this study was unavoidable. Even though we recruited three coders who manually reviewed the retrieved results by following reasonable criteria, the accuracy of the coding results was not fully under control. For instance, in Section 3.2, there was a “true” verdict, which should have been rated “irrelevant” in Section 2.2.1 and ignored because its claim assessed by fact-checkers (i.e., “Disposable masks should always be worn colored-side-out”) was not addressing the same issue as the original input claim (i.e., “Two ways of wearing a surgical mask: Colored side out if you are sick and white side out if you do not want to become sick”), that is, the verdict “true” did not indicate that the original claim was true but rather another irrelevant claim, but it was still rated “relevant” by two of the three coders and kept for data analysis thereafter. As for the evaluation of source reliability, even though we referenced two professional websites for media evaluation, as emphasized on their About pages, their ratings could also be more or less biased. Although subjectivity is unavoidable by nature in any qualitative research, future work may recruit more coders and leverage more rating sources and criteria to further reduce the potential influence of subjectivity. Finally, as we introduced in 2.1, even though fact-checking is applicable to any topic, this study centered on COVID-19-related misinformation due to its extensive rampancy, the severe consequence, and abundant datasets. Future studies may extend this study to other topics, such as politics, climate change, and influential social problems.